OpenStudy (anonymous):

lim x--> 3 x-3 / 2x^2 - 9x - 9

OpenStudy (anonymous):

wait

OpenStudy (freckles):

i would attempt to factor the denominator

OpenStudy (freckles):

then see if any factors cancel

OpenStudy (freckles):

or just plug in 3 and see if you need to do that

OpenStudy (freckles):

if you get 0/0 you need to if you don't then you are done

OpenStudy (anonymous):

Just plug in 3

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle\lim_{x \rightarrow ~3}~~\frac{x-3}{2x^2 - 9x - 9}}\)

OpenStudy (anonymous):

I am confused ... I do not understand what you are saying .. :(

OpenStudy (solomonzelman):

you have \(\large\color{black}{\displaystyle\lim_{x \rightarrow ~3}~~\frac{x-3}{2x^2 - 9x - 9}}\) and you are able to just plug in 3 for x, without doing anything, because when x=3, the denominator is NOT zero.

OpenStudy (solomonzelman):

can you tell me what you get after plugging in x=3?

OpenStudy (anonymous):

yes hold on

OpenStudy (solomonzelman):

sure

OpenStudy (anonymous):

do I plug in 3 to the top and bottom or only the denominator?

OpenStudy (solomonzelman):

you plug in three for x, into the top and the bottom.

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle\lim_{x \rightarrow ~3}~~\frac{x-3}{2x^2 - 9x - 9}=\frac{(3)-3}{2(3)^2 - 9(3) - 9}}\) like this

OpenStudy (anonymous):

i got 0

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

it is 0/-18 which is same as just 0.

OpenStudy (anonymous):

yeah

OpenStudy (solomonzelman):

good work!

OpenStudy (anonymous):

but that is all I do ? don't I have to get a number before and after 3 and plug it in to see what the limity as x approaches 3 is?

OpenStudy (anonymous):

u get a medal @SolomonZelman

OpenStudy (anonymous):

limit*

OpenStudy (solomonzelman):

I didn't do anything

OpenStudy (solomonzelman):

@ilovereyvis_x3 some other questions?

OpenStudy (solomonzelman):

btw do you want to know how I am putting up this in latex?

OpenStudy (anonymous):

my teacher doesn't do that , she never plugs in the number she finds a number before and after that number to find the limit as it approaches that certain number ... do you know what I am talking about?

OpenStudy (solomonzelman):

what does your teacher do, she factors or does other techniques to find the limit? (I mean in this case, since there is no restriction for x to equal 3. you CAN just plug in 3 for x. This is the correct way to do it.)

OpenStudy (freckles):

numerical approach?

OpenStudy (anonymous):

I am not sure what she calls it , it is weird , for example for this one we are finding the limit as x approaches 3 so she would use -2.9 and 3.1 or any number before or after the number 3

OpenStudy (solomonzelman):

ohh she is just filling in the table of values

OpenStudy (freckles):

that is called a numerical approach but usually more numbers are chosen to the left for left limit and usually more numbers are chosen to the right for right limit (you know instead of just one)

OpenStudy (solomonzelman):

lets say x is approaching 3. So she fills in values of that limit as x approach 3 from the left. values like x=2.5 x=2.9 x=2.99 x=2.999 plugging these values and seeing what the limit approaches from the left side. and from the right side she does. x=3.5 x=3.1 x=3.01 x=3.001

OpenStudy (solomonzelman):

like a big table of values

OpenStudy (anonymous):

I am not sure I am going to attach my notes so you can see a visual example of what I do in my class

OpenStudy (solomonzelman):

I know what approach you are taking about.

OpenStudy (solomonzelman):

see my big reply above?

OpenStudy (anonymous):

yes exactly like that !

OpenStudy (anonymous):

that is what she wants us to show as work

OpenStudy (solomonzelman):

yes, so we could have done it like if you want to. But once we have actually found the limit, it wouldn't make a difference what approach to use. We can still make a table if you want to.

OpenStudy (solomonzelman):

can you slow down please

OpenStudy (anonymous):

those are the notes

OpenStudy (solomonzelman):

yes, okay and what would you like to know?>

OpenStudy (anonymous):

how can I do the problem the way she wants it so I can understand it and get full credit because I have 6 problems to do with limits for this winter packet she gave me

OpenStudy (solomonzelman):

you have to estimate the limit. for example when x is approaching 4 of f(x) make 2 tables. left side right side \(\normalsize\color{ black }{\LARGE{\bbox[5pt, lightyellow ,border:2px solid white ]{ \huge{ \begin{array}{| l | c | r |} \hline \scr~~~x~~ & \scr f(x) \\ \hline \scr~~~3.5~ & \scr ~~~ \\ \hline \scr~~~3.7~ & \scr ~~~ \\ \hline \scr~~~3.9~ & \scr ~~~ \\ \hline \scr~~~3.99~ & \scr ~~ \\ \hline \scr~~~3.999~ & \scr ~~ \\ \hline \end{array} }}}}\) \(\normalsize\color{ black }{\LARGE{\bbox[5pt, lightyellow ,border:2px solid white ]{ \huge{ \begin{array}{| l | c | r |} \hline \scr~~~x~~ & \scr f(x) \\ \hline \scr~~~4.5~ & \scr ~~ \\ \hline \scr~~~4.3~ & \scr 3 \\ \hline \scr~~~4.1~ & \scr 4 \\ \hline \scr~~~4.01~ & \scr 5 \\ \hline \scr~~~4.001~ & \scr 5 \\ \hline \end{array} }}}}\)

OpenStudy (solomonzelman):

oops, they didn't fit. the bottom chart is for the right side, the top chart is for the left side.

OpenStudy (solomonzelman):

and any values of f(x) that are there disregard them

OpenStudy (solomonzelman):

See the Xs that I am plugging in?

OpenStudy (anonymous):

so I plug those numbers in as x ?

OpenStudy (solomonzelman):

not necessarily those. it is your choice what to plug in for x, unless you are given specific numbers to plug.

OpenStudy (anonymous):

Ok so I just need to pick one from the right and one from the left , which ones do you think will be best ?

OpenStudy (solomonzelman):

But you see that in both of my tables the x was getting closer and closer to 4. I could have made the charts smaller, apologize for thyat.

OpenStudy (anonymous):

for example if they say find the limit as x approaches 0 she would use 0.1 and -0.1

OpenStudy (anonymous):

yeah i see that

OpenStudy (anonymous):

so for this one I can use 2.9 and 3.5 ?

OpenStudy (solomonzelman):

yes, but the numbers in both of the tables have to be the same distance from 3 (if x is approaching 3)

OpenStudy (anonymous):

ok , so then which ones do you think I should use ?

OpenStudy (solomonzelman):

Considering a case where \(\large\color{black}{\displaystyle\lim_{x \rightarrow ~0}f(x)}\). (any f(x) ) left side right side \(\large\color{black}{ \large{ \begin{array}{| l | c | r |} \hline \scr~~~x~~ & \scr f(x) \\ \hline \scr-1 & \scr \\ \hline \scr-0.5 & \scr \\ \hline \scr-0.1 & \scr \\ \hline \scr-0.05 & \scr \\ \hline \scr-0.01 & \scr \\ \hline \scr-0.001 & \scr \\ \hline \end{array} } }\) \(\large\color{black}{ \large{ \begin{array}{| l | c | r |} \hline \scr~~~x~~ & \scr f(x) \\ \hline \scr1 & \scr \\ \hline \scr0.5 & \scr \\ \hline \scr0.1 & \scr \\ \hline \scr0.05 & \scr \\ \hline \scr0.01 & \scr \\ \hline \scr0.001 & \scr \\ \hline \end{array} } }\)

OpenStudy (solomonzelman):

I am not filling in the f(x) values, but just showing how I set up the X values for each side. they are away from the zero the same amount, the same distance. like -0.05 is just as far from 0 as 0.05 is.

OpenStudy (anonymous):

yeah for that case I use -0.1 and 0.1

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I see what your saying , so if I use 2.9 I would have to use -2.9

OpenStudy (anonymous):

I think ...

OpenStudy (solomonzelman):

but it it is Considering a case where \(\large\color{black}{\displaystyle\lim_{x \rightarrow ~2}f(x)}\). (any f(x) ) left side right side \(\large\color{black}{ \large{ \begin{array}{| l | c | r |} \hline \scr~~~x~~ & \scr f(x) \\ \hline \scr3 & \scr \\ \hline \scr2.5 & \scr \\ \hline \scr2.1 & \scr \\ \hline \scr2.05 & \scr \\ \hline \scr2.01 & \scr \\ \hline \scr2.001 & \scr \\ \hline \end{array} } }\) \(\large\color{black}{ \large{ \begin{array}{| l | c | r |} \hline \scr~~~x~~ & \scr f(x) \\ \hline \scr1 & \scr \\ \hline \scr1.5 & \scr \\ \hline \scr1.9 & \scr \\ \hline \scr1.95 & \scr \\ \hline \scr1.99 & \scr \\ \hline \scr1.999 & \scr \\ \hline \end{array} } }\)

OpenStudy (solomonzelman):

I mean the first table is rigth side and the second table is left side. sorry

OpenStudy (solomonzelman):

I have to leave for a second

OpenStudy (anonymous):

don't worry I got it , thank you for your help !

OpenStudy (solomonzelman):

you got it, sure? (I am back)